OpenStudy (anonymous):

Exponential to Logarithmic 4^3 = 64

OpenStudy (amistre64):

b^x = y logb(b^x) = logb(y) ; logb and b cancel each other and you are left with (x) x = logb(y) 4^3 = 64 log4(4^3) = log4(64) 3 = log4(64)

OpenStudy (amistre64):

A little more detail: logb(b^x) = x times logb(b); and logb(b) = 1